# Efficient Generation of Smooth Paths with Curvature Guarantees by Mollification

### v1

@@ Abstract
Most path following and trajectory tracking algorithms in mobile robotics require the desired path or trajectory to be defined by at least twice continuously differentiable functions to guarantee key properties such as global convergence, especially for nonholonomic robots like unicycles with speed constraints. Consequently, these algorithms typically exclude continuous but non-differentiable paths, such as piecewise functions. Despite this exclusion, such paths provide convenient high-level inputs for describing robot missions or behavior. While techniques such as spline interpolation or optimization-based methods are commonly used to smooth non-differentiable paths or create feasible ones from sequences of waypoints, they either can produce unnecessarily complex trajectories or are computationally expensive. In this work, we present a method to regularize non-differentiable functions and generate feasible paths through mollification. Specifically, we approximate an arbitrary path with a differentiable function that can converge to it with arbitrary precision. Additionally, we provide a systematic method for bounding the curvature of generated paths, which we demonstrate by applying it to paths resulting from linking a sequence of waypoints with segments. The proposed approach is computationally efficient, enabling real-time implementation on microcontrollers and compatibility with standard trajectory tracking and path following algorithms.
I Introduction
@@
Path generation transforms a set of points, e.g., a finite collection of waypoints, into a smooth, continuous curve that satisfies geometric properties such as continuity, bounded curvature, and differentiability, to respect vehicle kinematic constraints, ultimately transforming a sparse set of discrete locations into a followable geometric path. Spline-based methods, including cubic B-splines [1, 2] and Bézier curves [3], ensure $C^{2}$ continuity and bounded curvature but may produce high curvature and cumbersome paths when the spacing of waypoints is irregular [4]. Gaussian process regression [5] and kernel smoothing [6] generate $C^{\infty}$-smooth approximations, trading exact waypoint passage for improved smoothness at the cost of $O(n^{3})$ computational complexity and potential numerical instability. Optimization-based approaches [7, 8] incorporate explicit curvature bounds through convex formulations, enabling real-time computation but requiring careful parameter tuning [9]. Sampling-based methods, like RRT* [10] with post-smoothing offer probabilistic completeness but may yield suboptimal curvature or other metrics. In summary, it is not surprising that the chosen method balances curvature control, computational complexity, and application requirements for interpolation versus approximation.
@@
In this paper, we propose an inexpensive path generation method based on mollifying non-differentiable paths, e.g., piecewise functions, using mollifier functions. Mollifiers are smooth functions that, via convolution, approximate non-smooth functions arbitrarily close. They have been extensively used in partial differential equations [11] and in studying the non-vanishing of generalized Riemann zeta functions [12]. From an engineering perspective, mollifiers are used in signal and image processing to estimate the original probability distribution of variables under measurement error [13]. However, to the best of the authors’ knowledge, mollifiers have not been applied to path generation problems, despite their natural fit for robotics applications. For example, given a finite collection of ordered waypoints, it produces a path that can be arbitrarily close, in the Euclidean sense, to the segments connecting the waypoints sequentially. For this specific generated path, our method is well-suited for unicycle-like robots that travel at constant speed with heading-rate constraints; i.e., we can guarantee a maximum curvature for the generated path at the mild cost of slightly deviating from these connecting segments—an unavoidable trade-off, since the union of segments produces a non-differentiable path. Our approach is particularly appealing because it is computationally inexpensive as it can be run on small microcontrollers, and the resulting path, in its parametric representation, can be followed using path-following techniques based on modern guiding vector fields [14]. We also show how our method can generate 3D paths and guarantee several properties for generic sets of input points, including convexity preservation, enclosure of the generated path, and bounded maximum length when the input is also closed.
@@
The article is organized as follows, Section III introduces the mollifier functions and the problem formulation. Section IV explores which geometrical and analytical aspects of the original function are affected by mollification. Convexity, concavity and quasiconvexity are treated, as well as the length of the path and the curvature. Section VI validates the theoretical findings numerically and experimentally. We end the article with some conclusions in Section VII.
II Notation
@@
In this paper we use Lebesgue integration rather than Riemann integration to overcome some limitations with sets of measure zero and interchanging integral and limits. Therefore, we denote the Lebesgue measure in one dimension as $\lambda$, i.e., $\lambda:=\lambda_{1}$, and as $\lambda_{n}$ the $n$’th dimensional Lebesgue measure, and we consider the measure space $(X,\mathcal{B}(X),\lambda_{n})$ with $X\subseteq\mathbb{R}^{n}$, and $\mathcal{B}(X)$ is the Borel sigma-algebra of $X$; hence, when we say a function is measurable we are actually saying that it is Borel measurable. Consequently, all integrals in this paper must be thought as the Lebesgue integral with respect to the Lebesgue measure, even if the notation $\lambda$ is sometimes omitted in the integral.
For $p\in[1,\infty)$, we say that a measurable function $f$ is locally $p$-integrable, denoted $f\in L^{p}_{\mathrm{loc}}(X)$, if for each $x\in X$ there exists a neighborhood $U$ of $x$ such that $f|_{U}\in L^{p}(U)$.
We denote by $\operatorname{id}:X\to X$ the identity function defined as $\operatorname{id}(x)=x$. Then, the indicator function of a set $A$ is defined as
@@
For any two sets $X,Y$ we denote $C(X,Y)$ as the set of continuous functions from $X$ to $Y$ and for $n\in\mathbb{N}\cup\{\infty\}$ we denote as $C^{n}(X,Y)$ the set of $n$-times continuously differentiable functions from $X$ to $Y$. For a set $A$ we denote its closure as $\overline{A}$ and the support of a real-valued function is defined as $\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}}$, where $\operatorname{dom}f$ is the domain of $f$.
Finally, a (parametric) path in $\mathbb{R}^{n}$ is a function $f:X\subset\mathbb{R}\to\mathbb{R}^{n}$. Writing $f=(f_{1},\dots,f_{n})$, we call $f_{i}$ the $i$’th component of the path for $i\in\{1,\dots,n\}$.
III Path generation by mollification
III-A Path generation requirements
@@
Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a measurable path and $\{\varepsilon_{i}\}_{i=1}^{n}$ be a collection of positive real numbers. Find a new path $T_{\varepsilon}(f):\mathbb{R}\to\mathbb{R}^{n}$ where $T_{\varepsilon}(f)=:\{T_{\varepsilon_{i}}(f)\}_{i=1}^{n}$ and $T_{\varepsilon_{i}}$ is a functional that acts on each component of $f$ such that:
@@
    $T_{\varepsilon_{i}}(f_{i})\to f_{i}$ as $\varepsilon_{i}\to 0$ in some sense of convergence.
@@
    For $p\in\mathbb{N}$ with $p\geq 2$, $T_{\varepsilon_{i}}(f_{i})\in C^{p}(\mathbb{R},\mathbb{R})$ for $i\in\{1,\dots,n\}$.
@@
III-B Mollifiers for the path generation
@@
Theorem 1 ([11, Appendix C, Theorem 7])
@@
While it may seem otherwise, Theorem 1 does not solve Problem 1 entirely. Indeed, the second requirement is satisfied straightforwardly. Regarding the computational efficiency from the third requirement, the computation of $F$ is straightforward. Indeed, because of the compact and symmetric support of the mollifier due to Assumption 1, we have that $F_{\varepsilon}(x)=\int_{[-\varepsilon,\varepsilon]}f(x-t)\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)$ is an inexpensive numerical operation, and note the compact integration interval. Furthermore, note that $F_{i}^{(n)}:=(f_{i}*\varphi_{\varepsilon_{i}})^{(n)}=f_{i}*\varphi_{\varepsilon_{i}}^{(n)}$, and also note that computing $F_{i}^{(n)}(x)$ does not require the existence of $f_{i}^{(n)}(x)$.
@@
Finally, for the third statement, suppose now $f$ is continuous and take any compact set $K\subset\mathbb{R}$. Then $f$ is uniformly continuous on $K$. Thus, the same arguments as above can be followed noting that in this case $\delta$ depends on the compact set $K$. ∎
IV Key properties of the generated path
While Theorems 1 and 2 solve Problem 1, they do not provide further details about the properties of the resulting path. In this section, we characterize key properties of the generated path based on the input path. Specifically, we address questions such as: under what conditions does the output path preserve (local/quasi) convexity of the input? Does the output maintain monotonicity or other qualitative properties of the input, such as for step functions? How is the output path positioned relative to the input? How is the output path enclosed, and what is its length when the input is a closed path? We defer the analysis of the output path’s curvature to Section V, where we provide a detailed curvature analysis for the case when the input is a sequence of 2D or 3D waypoints. This curvature analysis serves as a systematic methodology that can be applied to other types of input paths.
@@ IV-A Convexity properties
Proposition 1 (Convexity and mollification)
@@
Proposition 2 (Local convexity and mollification)
@@
Note that by the selection of $\varepsilon$ and $\delta$ we have that $t\in(-\varepsilon,\varepsilon)\subset(-\delta,\delta)$. Thus, for any $t\in(-\varepsilon,\varepsilon)$, $\xi-t\in V$ and $\zeta-t\in V$. Since $V$ is convex we have that $\gamma(\xi-t)+(1-\gamma)(\zeta-t)\in V$ for all $t\in(-\varepsilon,\varepsilon)$ and $\gamma\in[0,1]$. Given that $f$ is convex in $(a,b)$ it is also convex in $V\subset(a,b)$, and noting $\varphi_{\varepsilon}\geq 0$ we can follow the steps of the proof of proposition 1 to reach
@@
Because $\xi,\zeta\in V$ and $\gamma\in[0,1]$ were arbitrary the proposition follows. ∎
Example 2
It is natural to ask whether Proposition 2 holds for any $\varepsilon>0$. That is, if $f$ is convex on an open set, is $F_{\varepsilon}$ also convex on that open set independently of $\varepsilon$ and the choice of non-negative mollifier? This is generally false, as demonstrated by the following counterexample. Consider the continuous function
@@
Nevertheless, for $\varepsilon=0.45$, we have that $F_{\varepsilon}$ is convex on a neighborhood contained in (0.5,0.5)(-0.5,0.5). Thus, we can set $\delta=0.45$ in Proposition 2, and for any $\varepsilon<\delta$, $F_{\varepsilon}$ is convex in (0.5,0.5)(-0.5,0.5).
Remark 1
Note that Propositions 1 and 2 also hold when $f$ is concave instead of convex, with concavity replacing convexity throughout.
@@
Proposition 3 (Affine functions and mollification)
@@
The result follows noting that $y\in\mathbb{R}\mapsto y\varphi_{\varepsilon}(y)$ is an odd function that is integrated over a symmetric interval. ∎
@@
Theorem 3 (Jensen’s inequality)
[15, Theorem 1.6.2] Let $\varphi$ be a non-negative measurable function such that $\int_{\mathbb{R}}\varphi\mathrm{d}\lambda=1$, $g$ be any measurable function and $f$ be a convex function such that $\operatorname{dom}f\supset\operatorname{img}g$; then
@@
Figure 3: Mollification of the function $f=|\operatorname{id}|$ with the mollifier $\varphi_{\varepsilon}$ as in Example 1 for different values of $\varepsilon$. Note that the function $f$ is convex, implying $F_{\varepsilon}$ is convex as demonstrated in proposition 1 and it is above the graph as shown in Proposition 4.
Now, we are ready to show that if $f$ is convex and $\varphi_{\varepsilon}$ is a non-negative even mollifier, then $\varphi_{\varepsilon}*f\geq f$ pointwise. For example, combined with Proposition 3, the following result implies that if our trajectory resembles an inverted tent, then the mollified function will also resemble a smoothed inverted tent but with the mollified function lying above the original path; see Figure 3 for the illustration.
Proposition 4 (If $f$ is convex then $F_{\varepsilon}$ is above $f$)
@@
The result now follows from Proposition 3 since $x$ is arbitrary. ∎
Remark 2
@@
The following result resembles Proposition 2, i.e., in order for the mollified function to be above the original function the parameter $\varepsilon$ has to be sufficiently small.
Proposition 5
@@
where the last equality comes from Proposition 3. Since $\xi\in(x,y)$ was arbitrary the result follows. ∎
Remark 3
Note that as it can be seen in Example 2, Proposition 5 does not hold for any $\varepsilon>0$. Indeed, also note that if the function is locally concave then we reach a similar results where the original function will be above the mollified function.
IV-B Quasiconvexity under mollification

### v2

# Efficient Path Generation with Curvature Guarantees by Mollification

@@ Abstract
Path generation, the process of converting high-level mission specifications, such as sequences of waypoints from a path planner, into smooth, executable paths, is a fundamental challenge in mobile robotics. Most path following and trajectory tracking algorithms require the desired path to be defined by at least twice continuously differentiable functions to guarantee key properties such as global convergence, especially for nonholonomic robots like unicycles with speed constraints. Consequently, path generation methods must bridge the gap between convenient but non-differentiable planning outputs, such as piecewise linear segments, and the differentiability requirements imposed by downstream control algorithms. While techniques such as spline interpolation or optimization-based methods are commonly used to smooth non-differentiable paths or create feasible ones from sequences of waypoints, they either produce unnecessarily complex trajectories or are computationally expensive. In this work, we present a method to regularize non-differentiable functions and generate feasible paths through mollification. Specifically, we approximate an arbitrary path with a differentiable function that can converge to it with arbitrary precision. Additionally, we provide a systematic method for bounding the curvature of generated paths, which we demonstrate by applying it to paths resulting from linking a sequence of waypoints with segments. The proposed approach is analytically shown to be computationally more efficient than standard interpolation methods, enabling real-time implementation on microcontrollers, while remaining compatible with standard trajectory tracking and path following algorithms.
1 Introduction
@@
In robotics and control, motion execution is often decomposed into a hierarchy of interrelated problems. Path planning determines a feasible route through the robot’s configuration space—typically as a sparse sequence of waypoints—using methods ranging from graph-search algorithms like A* to sampling-based approaches such as RRT [10]. Path generation then transforms this abstract plan into a smooth, continuous geometric path or time-parameterized trajectory, often using Spline-based methods, including cubic B-splines [9, 1] and Bézier curves [16]. These methods ensure $C^{2}$ continuity and bounded curvature but may produce high curvature and cumbersome paths when the spacing of waypoints is irregular [11]. Gaussian process regression [13] and kernel smoothing [5] generate $C^{\infty}$-smooth approximations, trading exact waypoint passage for improved smoothness at the cost of $O(n^{3})$ computational complexity and potential numerical instability. Optimization-based approaches [12, 14] incorporate explicit curvature bounds through convex formulations, enabling real-time computation but requiring careful parameter tuning [7].
However, critically, path generation does not operate in isolation: its output must meet the requirements imposed by the downstream control layer, which typically takes one of two forms. In path following, the generated path must be geometrically smooth and continuously differentiable, so that a well-defined tangent and curvature exist at every point for spatial error computation [17]. In trajectory tracking, the demands are more stringent: the time parameterization must additionally be consistent with the robot’s dynamic capabilities, as a reference trajectory that demands accelerations or velocities beyond the system’s limits cannot be reliably tracked regardless of the controller’s design [15]. In both cases, path generation must internalize the robot’s kinematic and dynamic constraints, effectively acting as the bridge that makes a planned route physically executable and controllable.
@@
In this paper, we propose an inexpensive path generation method based on mollifying non-differentiable paths, e.g., piecewise functions, using mollifier functions. Mollifiers are smooth functions that, via convolution, approximate non-smooth functions arbitrarily close. They have been extensively used in partial differential equations [4] and in studying the non-vanishing of generalized Riemann zeta functions [2]. From an engineering perspective, mollifiers are used in signal and image processing to estimate the original probability distribution of variables under measurement error [8]. However, to the best of the authors’ knowledge, mollifiers have not been applied to path generation problems, despite their natural fit for robotics applications.
Mollification offers a principled approach to path generation. For example, given a finite collection of ordered waypoints, it produces a smooth curve that approximates the piecewise-linear route to an arbitrary degree of Euclidean closeness. For this specific generated path, our method is well-suited for unicycle-like robots that travel at constant speed with heading-rate constraints; i.e., we can guarantee a maximum curvature for the generated path at the mild cost of slightly deviating from these connecting segments—an unavoidable trade-off, since the union of segments produces a non-differentiable path. Our approach is particularly appealing because it is computationally inexpensive as it can be run on small microcontrollers, and the resulting path, in its parametric representation, can be followed using path-following techniques based on modern guiding vector fields [17]. We also show how our method can generate 3D paths and guarantee several properties for generic sets of input points, including convexity preservation, enclosure of the generated path, and bounded maximum length when the input is also closed.
@@
The article is organized as follows, Section 3 introduces the mollifier functions and the requirements for the path generation problem. Section 4 explores which geometrical and analytical aspects of the original function are affected by mollification. Convexity, concavity and quasiconvexity are treated, as well as the length of the path and the curvature. Section 6 validates the theoretical findings numerically and experimentally. We end the article with some conclusions in Section 7.
2 Notation
@@
In this paper we use Lebesgue integration rather than Riemann integration to overcome some limitations with sets of measure zero and interchanging integral and limits. Therefore, we denote the Lebesgue measure in one dimension as $\lambda$, i.e., $\lambda:=\lambda_{1}$, and as $\lambda_{n}$ the $n$’th dimensional Lebesgue measure, and we consider that it is equipped with the Borel sigma algebra. Consequently, all integrals in this paper must be thought as the Lebesgue integral with respect to the Lebesgue measure, even if the notation $\lambda$ is sometimes omitted in the integral. For $p\in[1,\infty)$, we say that a measurable function $f$ is locally $p$-integrable, denoted $f\in L^{p}_{\mathrm{loc}}(X)$, if it is $p$-integrable in each compact subset of $X$. We denote by $\operatorname{id}:X\to X$ the identity function defined as $\operatorname{id}(x)=x$. Then, the indicator function of a set $A$ is defined as
@@
For any two sets $X,Y$ we denote $C(X,Y)$ as the set of continuous functions from $X$ to $Y$ and for $n\in\mathbb{N}\cup\{\infty\}$ we denote as $C^{n}(X,Y)$ the set of $n$-times continuously differentiable functions from $X$ to $Y$. For a set $A$ we denote its closure as $\overline{A}$ and the support of a real-valued function is defined as $\operatorname{supp}f:=\overline{\{x\in\operatorname{dom}f\mid f(x)\neq 0\}}$, where $\operatorname{dom}f$ is the domain of $f$. Finally, a (parametric) path in $\mathbb{R}^{n}$ is a measurable function $f:X\subset\mathbb{R}\to\mathbb{R}^{n}$. Writing $f=(f_{1},\dots,f_{n})$, we call $f_{i}$ the $i$’th component of the path for $i\in\{1,\dots,n\}$.
3 Path generation by mollification
3.1 Path generation requirements
@@
Let $f:\mathbb{R}\to\mathbb{R}^{n}$ be a parametric path and $\{\varepsilon_{i}\}_{i=1}^{n}$ be a collection of positive real numbers. Find a new path $T_{\varepsilon}(f):\mathbb{R}\to\mathbb{R}^{n}$ where $T_{\varepsilon}(f)=:\{T_{\varepsilon_{i}}(f_{i})\}_{i=1}^{n}$ and $T_{\varepsilon_{i}}$ is a functional that acts on each component of $f$ such that:
@@
    Each component of the parametric path can be made arbitrary close to the original path, that is, $T_{\varepsilon_{i}}(f_{i})\to f_{i}$ as $\varepsilon_{i}\to 0$ in some sense of convergence.
@@
    It provides enough smoothness, that is, for $p\in\mathbb{N}$ with $p\geq 2$, $T_{\varepsilon_{i}}(f_{i})\in C^{p}(\mathbb{R},\mathbb{R})$ for $i\in\{1,\dots,n\}$.
@@
3.2 Mollifiers for the path generation
@@
Theorem 1 ([4, Appendix C, Theorem 7])
@@
While it may seem otherwise, Theorem 1 does not solve Problem 1 entirely. Indeed, the second requirement is satisfied straightforwardly. Regarding the computational efficiency from the third requirement, the computation of $F$ is also straightforward. Indeed, because of the compact and symmetric support of the mollifier due to Assumption 1, we have that $F_{\varepsilon}(x)=\int_{[-\varepsilon,\varepsilon]}f(x-t)\varphi_{\varepsilon}(t)\,\mathrm{d}\lambda(t)$ is an inexpensive numerical operation, and note the compact integration interval. Furthermore, note that $F_{i}^{(n)}:=(f_{i}*\varphi_{\varepsilon_{i}})^{(n)}=f_{i}*\varphi_{\varepsilon_{i}}^{(n)}$, and also note that computing $F_{i}^{(n)}(x)$ does not require the existence of $f_{i}^{(n)}(x)$.
@@
Finally, for the third statement, suppose now $f$ is continuous and take any compact set $K\subset\mathbb{R}$. Then $f$ is uniformly continuous on $K$. Thus, the same arguments as above can be followed noting that in this case $\delta$ depends on the compact set $K$.
4 Key properties of the generated path
While Theorems 1 and 2 solve Problem 1, they do not provide further details about the properties of the resulting path. In this section, we characterize key properties of the generated path based on the input path. Specifically, we address questions such as: under what conditions does the output path preserve (local/quasi) convexity of the input? Does the output maintain monotonicity or other qualitative properties of the input, such as for step functions? How is the output path positioned relative to the input? How is the output path enclosed, and what is its length when the input is a closed path? We defer the analysis of the output path’s curvature to Section 5, where we provide a detailed curvature analysis for the case when the input is a sequence of 2D or 3D waypoints. This curvature analysis serves as a systematic methodology that can be applied to other types of input paths.
@@ 4.1 Convexity properties
Proposition 4.4 (Convexity and mollification).
@@
Proposition 4.6 (Local convexity and mollification).
@@
Note that by the selection of $\varepsilon$ and $\delta$ we have that $t\in(-\varepsilon,\varepsilon)\subset(-\delta,\delta)$. Thus, for any $t\in(-\varepsilon,\varepsilon)$, $\xi-t\in V$ and $\zeta-t\in V$. Since $V$ is convex we have that $\gamma(\xi-t)+(1-\gamma)(\zeta-t)\in V$ for all $t\in(-\varepsilon,\varepsilon)$ and $\gamma\in[0,1]$. Given that $f$ is convex in $(a,b)$ it is also convex in $V\subset(a,b)$, and noting $\varphi_{\varepsilon}\geq 0$ we can follow the steps of the proof of proposition 4.4 to reach
@@
Because $\xi,\zeta\in V$ and $\gamma\in[0,1]$ were arbitrary the proposition follows.
Example 4.8.
It is natural to ask whether Proposition 4.6 holds for any $\varepsilon>0$. That is, if $f$ is convex on an open set, is $F_{\varepsilon}$ also convex on that open set independently of $\varepsilon$ and the choice of non-negative mollifier? This is generally false, as demonstrated by the following counterexample. Consider the continuous function
@@
Nevertheless, for $\varepsilon=0.45$, we have that $F_{\varepsilon}$ is convex on a neighborhood contained in (0.5,0.5)(-0.5,0.5). Thus, we can set $\delta=0.45$ in Proposition 4.6, and for any $\varepsilon<\delta$, $F_{\varepsilon}$ is convex in (0.5,0.5)(-0.5,0.5).
Remark 4.9.
Note that Propositions 4.4 and 4.6 also hold when $f$ is concave instead of convex, with concavity replacing convexity throughout.
@@
Proposition 4.10 (Affine functions and mollification).
@@
The result follows noting that $y\in\mathbb{R}\mapsto y\varphi_{\varepsilon}(y)$ is an odd function that is integrated over a symmetric interval.
@@
Theorem 4.12 (Jensen’s inequality).
[3, Theorem 1.6.2] Let $\varphi$ be a non-negative measurable function such that $\int_{\mathbb{R}}\varphi\mathrm{d}\lambda=1$, $g$ be any measurable function and $f$ be a convex function such that $\operatorname{dom}f\supset\operatorname{img}g$; then
@@
Figure 3: Mollification of the function $f=|\operatorname{id}|$ with the mollifier $\varphi_{\varepsilon}$ as in Example 1 for different values of $\varepsilon$. Note that the function $f$ is convex, implying $F_{\varepsilon}$ is convex as demonstrated in proposition 4.4 and it is above the graph as shown in Proposition 4.13.
Now, we are ready to show that if $f$ is convex and $\varphi_{\varepsilon}$ is a non-negative even mollifier, then $\varphi_{\varepsilon}*f\geq f$ pointwise. For example, combined with Proposition 4.10, the following result implies that if our trajectory resembles an inverted tent, then the mollified function will also resemble a smoothed inverted tent but with the mollified function lying above the original path; see Figure 3 for the illustration.
Proposition 4.13 (If $f$ is convex then $F_{\varepsilon}$ is above $f$).
@@
The result now follows from Proposition 4.10 since $x$ is arbitrary.
Remark 4.15.
@@
The following result resembles Proposition 4.6.
Proposition 4.16 (Local convexity and $F_{\varepsilon}\geq f$).
@@
where the last equality comes from Proposition 4.10. Since $\xi\in(x,y)$ was arbitrary the result follows.
Remark 4.18.
Note that as it can be seen in Example 4.8, Proposition 4.16 does not hold for any $\varepsilon>0$. Indeed, also note that if the function is locally concave then we reach a similar result where the original function will be above the mollified function.
4.2 Quasiconvexity under mollification